# Localising the Seizure Onset Zone from Single-Pulse Electrical Stimulation Responses with a Transformer

### v1

@@ 5.1 Performance metrics
Table 1 presents the performance metrics for each model. CNNdivergent, previously shown to attain a Youden’s index of 0.527 (Johnson et al., 2022), underperformed in this task, registering a Youden’s index of only 0.098 and an AUROC marginally above chance at 0.574. In comparison, CNNconvergent had an increased AUROC of 0.666, yet it still faced challenges in generalisation.
We observed improved model performance in Transformerbase, which attained an AUROC of 0.730. This improvement over CNNconvergent was statistically significant ($p=0.036$). Incorporating additional features in Transformerall yielded a minor increase in the AUROC to 0.745, though the AUPRC remained almost constant (0.377 to 0.378). While this suggests only a marginal improvement at the individual patient level – if any – a notable increase in the Youden index from 0.233 to 0.305 was observed, mainly due to a large increase in sensitivity (0.597 to 0.663). This improvement, achieved using a threshold determined from the validation set, points to an enhanced ability of the model to generalise across different patients.
ROC curves for each model are shown in Appendix F.
@@ 5.2 Channel sensitivity analysis
The channel sensitivity analysis assessed the relationship between Transformerbase performance and the number of channels used during inference. A crucial finding was the model’s robust performance across varying channel counts, including a monotonic increase in performance with an increasing number of channels, as illustrated in Figure LABEL:fig:ablation.
When the channel count was set to 37, in line with CNNconvergent, our model achieved an AUROC of 0.724. This performance notably surpassed that of the CNN model, which had an AUROC of 0.666 with the same number of channels. Furthermore, the resilience of Transformerbase was underscored when the channel count was reduced to just 4, yielding an AUROC of 0.667, equalling that of CNNconvergent.
@@
Figure 3: Channel sensitivity analysis for Transformerbase. In this analysis, a random subset of $n$ channels served as input for each prediction, testing different values of $n$. This analysis used the previously trained Transformerbase models. AUROC scores for CNNconvergent, which used 37 channels, are presented for comparison. Bootstrapping was used to calculate 95% confidence intervals.
[…]
fig:ft_nft_scores Refer to caption
Figure 4: Frontal/temporal lobe SOZ analysis. The trained Transformer models were evaluated on two subcohorts: those with a majority of frontal/temporal lobe SOZs and those with a minority. For each of Transformerbase and Transformerall, AUROC scores were calculated across these subcohorts, with 95% confidence intervals calculated using bootstrapping.
5.3 Transformer performance in Frontal/Temporal SOZs
This analysis focused on assessing whether the improved performance of Transformerbase was more pronounced in the cohort with predominantly frontal/temporal lobe SOZs, which could suggest the identification of delayed responses. However, the data did not support an improvement in AUROC scores within this cohort (0.741 to 0.743). Instead, we observed an overall increase in the mean AUROC in the cohort with a minority of frontal/temporal lobe SOZs (0.702 to 0.750), albeit with overlapping confidence intervals. Figure LABEL:fig:ft_nft_scores illustrates these findings.
@@ 5.4 Predictions visualised
Example predictions for a single patient are visualised in Figure LABEL:fig:predictions. CNNconvergent significantly overlaps with the clinician labelled SOZ, albeit with several false positives. Transformerbase exhibits a similar overlap but with fewer false positives. Notably, Transformerall demonstrates the most effective performance, with substantial alignment with the clinician labels, reflecting the superior capabilities of this model.
@@
Figure 5: Model outputs for a single patient. Electrodes are classified as SOZ or non-SOZ. Nodes represent electrodes; colours denote clinician-determined ground truth (top left) or the prediction for a given model. This patient was absent from the training set, demonstrating model generalisability to unseen patients and electrodes.
@@ 6.1 Baseline performance
The heterogeneity of seizure onsets in this dataset (see Appendix A) compared with the previous focus on temporal lobe epilepsies may have been a contributing factor. Given the limited sample size, this may have led to underfitting. Another possible explanation is that our approach averaged responses across trials, attenuating delayed responses. However, Johnson et al. (2022) reported only a slight decrease in performance when limiting the post-stimulation window to 0-175ms – i.e., removing most delayed responses, suggesting any performance reduction attributable to averaging across trials was likely negligible.
@@ 6.2 Divergent vs Convergent paradigms
In the introduction, we hypothesised that a convergent paradigm might be more conducive to effective SOZ localisation. This hypothesis was premised on the literature suggesting that response sites are more indicative of epileptogenicity than stimulation sites. Our empirical investigation, comparing CNNdivergent with CNNconvergent, lends credence to this hypothesis. The AUROC increased notably from 0.574 to 0.666, marking a substantial improvement, although this remains low for a task of critical importance.
6.3 CNN vs Transformer comparison
@@
In our comparative study of CNNconvergent and Transformerbase, the latter demonstrated statistically significant superiority. An important question was whether this advantage was solely due to its ability to process all channels during inference rather than just a subset. Section 5.2 shows evidence to the contrary, where Transformerbase outperformed CNNconvergent even when restricted to the same number of inference channels.
@@
The Transformer’s cross-channel attention mechanism enables improved feature extraction compared to CNNconvergent, which applies a ResNet to compute a single set of features across channels. The superiority of the Transformer lies in its ability to extract channel-wise features and then apply cross-channel attention, explaining the performance disparity between the two models. Recent evidence has demonstrated the efficacy of cross-channel attention in intracranial EEG analysis (Zhang et al., 2024). Our findings further extend its applicability, demonstrating its ability to model evoked responses effectively. Given the heterogeneity in subject-specific electrode configurations and its high spatial resolution, intracranial EEG is particularly well-suited to benefit from this mechanism.
6.4 Additional Transformer features
The lack of improvement in the frontal/temporal lobe SOZ cohort for Transformerall suggests a lack of improvement in detecting delayed responses. Had this been the case, we would expect an increase in performance in this cohort, given that delayed responses are seen in these regions (Valentín et al., 2005). Unless Transformerbase is already adept at detecting delayed responses, this indicates the potential need for a more sophisticated approach to detect such responses. For instance, a method that explicitly considers all trials might be required instead of relying on summary statistics across trials.
It is worth considering why AUROC scores for the remainder of the cohort may have improved. Firstly, the standard deviation may reflect other forms of inter-trial variability, which Cornblath et al. (2023) suggests could possess predictive power in identifying epileptogenic tissue. Secondly, as discussed in Section 3.2.4, the addition of an extra branch in the per-channel embedding might have enhanced the discrimination between epileptogenic and non-epileptogenic CCEPs. Finally, incorporating the distance between stimulating and recording electrodes as a feature may have contributed to this increased generalisability.
6.5 Limitations and future work
A fundamental limitation of our models is that even the best-performing method, Transformerconvergent, only has a Youden score of 0.305. Without considering patient outcomes, it is difficult to know the true capability of these methods, as patients with poor outcomes would likely have a significant impact on model performance. Future work should aim to train models on patients with favourable outcomes and compare the performance between unseen patients with bad and good outcomes. Improved performance on patients with favourable outcomes would indicate that these models are more indicative of epileptogenic regions than SOZ labels alone.
@@ 6.5 Limitations and future work
Another challenge is the limited scope of our models’ validation, which is confined to patients from the same dataset. Whilst we evaluated generalisability to unseen patients and electrode placements, our evaluation did not test for generalisability to other centres, a fundamental limitation given the variability in stimulation protocols and electrode modality (ECoG or SEEG). To validate their generalisability convincingly, testing on an external dataset is paramount. Ideally, such testing should be preceded by training models on a dataset derived from a multi-centre cohort, thus ensuring model robustness across clinical environments.
Aligning our strategy with clinical practice is another area for development. Clinical practice incorporates diverse data, including seizure activity, interictal recordings, magnetic resonance and PET imaging, and other clinical information. This integration takes place during multidisciplinary team meetings through discussions and visual comparisons. A multi-modal machine learning approach, utilising extensive, high-quality data, could significantly improve model performance, assuming complexity and scalability issues are overcome (Kline et al., 2022).
@@
Finally, as discussed in section 6.4, our approach of using summary statistics across trials is likely suboptimal. While this method captures some inter-trial variability whilst increasing the signal-to-noise ratio of CCEPs, it may only partially capture trial-by-trial variability, obscuring trends over time. Learning directly across trials could provide a more nuanced understanding of these responses.
@@ 7 Conclusion
This paper presents several contributions for modelling intracranial EEG data using deep learning, particularly in localising the SOZ from SPES data. Our research first compares divergent and convergent paradigms in classifying multi-channel SPES responses. This comparison revealed a substantial improvement in moving from the conventional divergent approach (AUROC: 0.574) to a convergent one (AUROC: 0.666).
Moreover, we have pioneered the use of Transformer models in analysing SPES responses. Our approach of applying a Transformer with cross-channel attention to invasive EEG data has shown promising results, elevating the AUROC to 0.730. Further refinement of this model by incorporating features such as inter-trial variability led to an increased AUROC of 0.745, demonstrating more consistent predictions across patients.
We aim to extend our work in several directions. One key area is investigating whether our approach can predict epileptogenicity more accurately than traditional SOZ localisation by considering patient outcomes. Another avenue is to explore the potential of the Transformer model to process SPES responses across trials rather than using summary statistics. Additionally, evaluating our model on an external dataset will be crucial, given the heterogeneity in SPES protocols.
Our research marks a notable advancement in modelling subject-specific, heterogeneous intracranial EEG placements in SPES. Through the application of deep learning methodologies, we strive to make a meaningful contribution towards enhancing surgical outcomes for individuals with drug-resistant epilepsy.
Appendix A Dataset
The processed dataset comprises 2066 electrodes, averaging 59 per patient. These were restricted to stimulated electrodes to facilitate a fair comparison between both paradigms. Among these electrodes, 298 (14.4%) were identified as within the SOZ. The electrode location distribution, including those within the SOZ, was consistent across lobes (Figure LABEL:fig:lobe_distribution), indicating that brain regions are unlikely to significantly confound the results.
@@ Appendix A Dataset
fig:lobe_distribution Refer to caption
Appendix B Dataset split
@@ Appendix C Preprocessing
Appendix D Metrics
The AUROC and AUPRC scores consider class imbalance, with AUPRC being particularly effective in imbalanced dataset scenarios. These metrics indicate, at the patient level, whether SOZ electrodes are generally scored higher than non-SOZ electrodes.
A threshold was required to compute the specificity, sensitivity, and Youden’s index. For each model, the optimal threshold was determined on the validation set to maximise Youden’s index and then applied to the corresponding test set. This approach assesses the practicality of using a consistent, patient-independent threshold and evaluates if the model’s ability to distinguish between SOZ and non-SOZ channels holds across different patients.
The statistical test between CNNconvergent and Transformerbase was a one-sided t-test using Nadeau and Bengio (1999)’s corrected method for repeated k-fold cross-validation. We calculated the p-value based on the mean AUROC scores from each fold and run, with a significance threshold set at 0.05.
Appendix E Model training procedures
@@
Ten trials were conducted for each model, a decision aimed at balancing model performance against the constraints of computational resources. Training was halted after ten epochs, and the set of hyperparameters that maximised the validation AUROC were selected. This process was only done once, with a separate random seed to the 5 runs, as this process took several times longer than the final model training. The best hyperparameters for the CNN and Transformer models are shown in Tables 2 and 3, respectively.
@@
Table 3: Hyperparameters for Transformer Models. The table presents the optimal dropout rate, embedding dimension, and number of layers for both Transformer models as determined by the hyperparameter search.
@@
Transformerbase 0.46 16 2 1.5 ×$\times 10^{-4}$
Transformerall 0.44 16 2 3.4 ×$\times 10^{-3}$
Appendix F ROC curves
Each model’s Receiver Operating Characteristic (ROC) curves are shown in Figure LABEL:fig:roc_curves, summarising the trade-off between sensitivity and specificity across various threshold settings. Consistent with the AUROC scores, the convergent paradigm outperformed the divergent paradigm across thresholds, and the Transformer models demonstrated superior performance compared to the CNN models. However, the ROC curves for the Transformer models displayed a notable similarity. In summary, these curves illustrate the superior discrimination ability of the Transformer models.
[…]
fig:roc_curves Refer to caption
Figure 7: ROC Curves for each model. Receiver Operating Characteristic (ROC) curves are shown for each model alongside a baseline random classifier. The AUROC scores are indicated for each model.

### v2

@@ 5.1 Performance metrics
[…]
fig:roc_curve_boxplot Refer to caption
Figure 4: A. ROC Curves. Receiver Operating Characteristic (ROC) curves are shown for each model alongside a baseline random classifier. The AUROC scores are indicated for each model. B. CNN Transformer performance by surgical outcome. The per-patient AUROC scores were stratified by surgical outcome for the CNN Transformer. Scores were generally higher in the seizure-free cohort (median: 0.86) than in the non-seizure-free cohort (median: 0.77), suggesting that the models may predict outcomes. However, this difference was not statistically significant.
Table 1 presents the performance metrics for each model. CNNdivergent, which was previously shown to attain a Youden’s index of 0.527 (Johnson et al., 2022), underperformed in this task, registering a Youden’s index of only 0.109 and an AUROC marginally above chance at 0.574. In comparison, CNNconvergent had an increased AUROC of 0.666, yet it still faced challenges in generalisation. When using cross-channel attention, the AUROC was 0.730. This improvement over CNNconvergent was statistically significant at the 5% level ($p=0.036$).
After selecting an optimal threshold, this hierarchy of model performance was maintained. The specificity and sensitivity values monotonically increased when switching to a convergent paradigm and then to a CNN Transformer. As the AUROC scores suggest, this hierarchy of model performance was consistent across thresholds (Figure LABEL:fig:roc_curve_boxplotA).
Figure LABEL:fig:roc_curve_boxplotB shows the distribution of AUROC scores for each outcome type. The general trend is that AUROC scores are higher in seizure-free patients following surgery ($n=11$), with a median AUROC of 0.86, compared to 0.77 in those not seizure-free ($n=23$). Furthermore, the lowest AUROC score was in a non-seizure-free patient. However, any improvement in AUROC was not statistically significant at the 5% level ($p=0.115$).
@@ 5.2 Channel sensitivity analysis
The channel sensitivity analysis assessed the relationship between the performance of the CNN Transformer and the number of channels used during inference. A crucial finding was the model’s robust performance across varying channel counts, including a monotonic increase in performance with an increasing number of channels, as illustrated in Figure LABEL:fig:ablation.
When the channel count was set to 37, in line with CNNconvergent, our model achieved an AUROC of 0.724. This performance notably surpassed that of the CNN model, which had an AUROC of 0.666 with the same number of channels. Furthermore, the resilience of the CNN Transformer was underscored when the channel count was reduced to just 4, with an AUROC of 0.667, equalling that of CNNconvergent.
@@
Figure 5: Channel sensitivity analysis for CNN Transformer. In this analysis, a random subset of $n$ channels served as input for each prediction, testing different values of $n$. This analysis used the previously trained CNN Transformer models. AUROC scores for CNNconvergent, which used 37 channels, are presented for comparison. Bootstrapping was used to calculate 95% confidence intervals.
@@ 5.3 Predictions visualised
Example predictions for a single patient are visualised in Figure LABEL:fig:predictions. Both CNNdivergent and CNNconvergent had a significant overlap with the clinician labelled SOZ, albeit with several false positives. The CNN Transformer exhibits a similar overlap but with fewer false positives.
@@
Figure 6: Model outputs for a single patient. (Left) Clinician-determined SOZ for an example patient. Nodes represent intracranial EEG electrodes. (Rest) Classification outputs for each model. Predictions from the CNN Transformer are most closely aligned with the clinician-determined SOZ. This patient was absent from the training set, demonstrating model generalisability to unseen patients and electrodes.
@@ 6.1 Baseline performance
The heterogeneity of seizure onsets in this dataset (see section 4.1) compared with the previous focus on temporal lobe epilepsies may have been a contributing factor. Given the limited sample size, this may have led to underfitting. Another possible explanation is that our approach averaged responses across trials, attenuating delayed responses. However, Johnson et al. (2022) reported only a slight decrease in performance when limiting the post-stimulation window to 0-175ms – i.e., removing most delayed responses, suggesting any performance reduction attributable to averaging across trials was likely negligible.
@@ 6.2 Divergent vs Convergent paradigms
In the introduction, we hypothesised that a convergent paradigm might be more conducive to effective SOZ localisation. This hypothesis was premised on the literature suggesting that response sites are more indicative of epileptogenicity than stimulation sites. Our empirical investigation, which compared CNNdivergent with CNNconvergent, supports this hypothesis. The AUROC increased notably from 0.574 to 0.666, marking a substantial improvement, although this remains low for a task of critical importance.
6.3 CNN vs CNN Transformer comparison
@@
In our comparative study of CNNconvergent and the CNN Transformer, the latter demonstrated statistically significant superiority. An important question was whether this advantage was solely due to its ability to process all channels during inference rather than just a subset. Section 5.2 shows evidence to the contrary, where the CNN Transformer outperformed CNNconvergent even when restricted to the same number of inference channels.
@@
The CNN Transformer’s cross-channel attention mechanism enables improved feature extraction compared to CNNconvergent, which applies a ResNet to compute a single set of features across channels. The superiority of the Transformer lies in its ability to extract channel-wise features and then apply cross-channel attention, explaining the performance disparity between the two models. Recent evidence has demonstrated the efficacy of cross-channel attention in intracranial EEG analysis (Zhang et al., 2024). Our findings further extend its applicability, demonstrating its ability to model evoked responses effectively. Given the heterogeneity in patient-specific electrode configurations and its high spatial resolution, intracranial EEG is particularly well-suited to benefit from this mechanism.
6.4 Model performance by outcome
While the increase in AUROC in the seizure-free cohort may not have reached statistical significance, the emerging trend suggests that there is potential for a significant increase, given the right conditions:
  1. 1.
    The small sample size ($n=34$) could be a limiting factor, as a larger cohort size is likely necessary to observe a significant increase in AUROC.
  2. 2.
    The models were trained using data from patients with all outcomes. For patients who were not rendered seizure-free after surgery, some SOZ channels may not have been epileptogenic, and some non-SOZ channels may have been epileptogenic.
  3. 3.
    The SOZ and resected channels typically diverge. A poor outcome does not imply that the SOZ did not indicate epileptogenicity, as removal of the SOZ may not have been possible.
Whilst the overall Youden score of the CNN Transformer is only 0.319, which is likely some way off being clinically useful, the performance in the seizure-free cohort – despite being trained on patients with all outcomes – suggests the potential for improved clinical utility. A move towards predicting surgical outcomes is one such approach, in which the training data would be restricted to patients who went on to be seizure-free. Channels that were both SOZ and resected would be labelled as positive instances (epileptogenic), and those that were neither SOZ nor resected as negative instances (non-epileptogenic). We could then investigate whether resecting more of the highest scoring channels – as determined by our models – relates to surgical outcome.
6.5 Clinical applications
Our current models demonstrate significant potential to enhance clinical practice in epilepsy treatment by indicating the likelihood of an epileptogenic focus.
Firstly, we propose presenting the model’s outputs visually within a 3D spatial context, enabling clinicians to intuitively visualise abnormal brain regions. This visualisation can be integrated with multimodal pre-surgical evaluation data, such as radiological findings, seizure activity, interictal recordings, and PET imaging, to align with current clinical practice. In future, a multi-modal machine learning approach, utilising extensive, high-quality data, could significantly improve model performance, assuming complexity and scalability issues are overcome (Kline et al., 2022).
Secondly, our model can flag areas where clinical labels and model predictions diverge, such as SOZ regions with low model scores or non-SOZ regions with high scores. Given the subjectivity in SOZ labelling, this feature can help clinicians identify additional regions of interest that may warrant further investigation, thus enhancing evaluation thoroughness beyond visual intracranial EEG analysis alone. Thirdly, providing continuous scores for each channel allows for a more nuanced assessment of epileptogenic activity likelihood. This detailed information supports clinicians in prioritising regions for surgical intervention and making more informed patient management decisions.
An alternative approach involves predicting surgical outcomes by restricting the training data to patients who became seizure-free post-surgery. Channels labelled as SOZ and resected would be positive instances (epileptogenic), while those neither SOZ nor resected would be negative instances (non-epileptogenic). This approach could help investigate whether resecting more of the highest-scoring channels, as determined by our models, correlates with better surgical outcomes. Further validation is crucial to confirm the effectiveness of these models in diverse patient populations and clinical settings, as discussed in subsequent sections.
Limitations
@@ 6.5 Clinical applications
A key challenge is the limited scope of our models’ validation, which was limited to patients from the same dataset. Whilst we evaluated generalisability to unseen patients and electrode placements, our evaluation did not test for generalisability to other centres, which is particularly important given the variability in stimulation protocols and electrode modality (ECoG or SEEG). To validate their generalisability convincingly, testing on an external dataset is paramount. Ideally, such testing should be preceded by training models on a dataset derived from a multi-centre cohort, thus ensuring model robustness across clinical environments. Furthermore, we did not conduct a thorough analysis of the relationship between model performance and factors such as age, SOZ location, and electrode placement. Future validation studies on larger cohorts will be necessary to account for these potential effects.
A further divergence from clinical practice lies in the models’ independently classifying each site. Intuitively, if one electrode site is deemed epileptogenic, we may also expect adjacent regions to be epileptogenic. Nonetheless, our current models do not account for this, as they make predictions for each electrode separately. Implementing post-processing techniques that consider the spatial relationship between electrodes could be beneficial in addressing this limitation.
@@
Finally, as discussed in section 1, various forms of inter-trial variability can indicate epileptogenicity. Averaging across trials increased the signal-to-noise ratio of SPES responses at the expense of these markers. Several approaches can be used to factor in this variability, such as capturing the standard deviation across trials. Alternatively, cross-trial attention can be used to directly extract features from relevant trials. This approach may be particularly suited to detecting delayed responses, which predict surgical outcome (Valentín et al., 2005).
@@ 7 Conclusion
This paper presents several contributions for modelling intracranial EEG data using deep learning, particularly in localising the SOZ from SPES data. Our research first compares divergent and convergent paradigms in classifying multi-channel SPES responses. This comparison revealed a substantial improvement in moving from the conventional divergent approach (AUROC: 0.574) to a convergent one (AUROC: 0.666). Our approach of applying a Transformer with cross-channel attention to invasive EEG data has shown promising results, elevating the AUROC to 0.730.
We aim to extend our work in several directions. One key area is investigating whether our approach can predict epileptogenicity more accurately than traditional SOZ localisation. There is currently only limited evidence for this, as any improvement in performance in the seizure-free cohort did not reach statistical significance. Another avenue is to explore the potential of the CNN Transformer to process SPES responses across trials rather than taking the mean. Additionally, evaluating our model on an external dataset will be crucial, given the heterogeneity in SPES protocols.
Our study significantly advances the modelling of heterogeneous intracranial EEG placements in SPES. We hope that further developments can improve surgical outcomes for patients with drug-resistant epilepsy.
@@
JN is supported by a UCL UKRI Centre for Doctoral Training in AI-enabled Healthcare studentship (EP/S021612/1).
Appendix A Dataset split
@@ Appendix B Preprocessing
Appendix C Metrics
The AUROC score considers class imbalance, with a score of 0.5 corresponding to a random classifier and 1.0 to a perfect classifier. This metric indicates, at the patient level, whether SOZ electrodes are generally scored higher than non-SOZ electrodes.
A threshold was required to compute the specificity, sensitivity, and Youden’s index. We determined patient-specific thresholds in the following manner:
  1. 1.
    For each patient in the validation set, we collated model predictions (SOZ likelihood) for each electrode ($x$).
  2. 2.
    We determined the patient-specific threshold that maximised the Youden score ($thresh$).
  3. 3.
    We calculated $n$ which satisfied the following: $thresh=median(x)+n\times MAD(x)$, where $MAD$ is the median absolute deviation.
  4. 4.
    We calculated the median $n$ across patients in the validation set.
This allowed us to calculate a patient-specific threshold for each patient in the test set that required no knowledge of which channels were SOZ.
The statistical test between CNNconvergent and the CNN Transformer was a one-sided t-test using Nadeau and Bengio (1999)’s corrected method for repeated k-fold cross-validation. We calculated the p-value based on the mean AUROC scores from each fold and run, with a significance threshold set at 0.05.
Appendix D Model training procedures
@@
Ten trials were conducted for each model, a decision aimed at balancing model performance against the constraints of computational resources. Training was halted after ten epochs, and the set of hyperparameters that maximised the validation AUROC were selected. This process was only done once, with a separate random seed to the 5 runs, as this process took several times longer than the final model training. The best hyperparameters for the CNN and CNN Transformer models are shown in Tables 2 and 3, respectively.
@@
Table 3: Hyperparameters for the CNN Transformer. The table presents the optimal dropout rate, embedding dimension, and number of layers for the CNN Transformer.
@@
CNN Transformer 0.46 16 2 1.5 ×$\times 10^{-4}$